[Hallc_running] [New Logentry] Hall C RC Update: Tuesday, Dec. 21, 2021 (last RC update of 2021)

gaskelld at jlab.org gaskelld at jlab.org
Tue Dec 21 08:45:01 EST 2021

Logentry Text:
After recovery from the issues with 1L05 yesterday, we continued taking data on LH2 at the SHMS left (Q2=6, W=2.4, low epsilon) setting.  We stayed on this setting until beam went off this morning, accumulating 19,873 good pion events.  Yesterday afternoon, we also took a BCM calibration (5 to 70 uA).

Beam went off this morning a few minutes after 7 am for the holiday shutdown.  Hall C is now in Restricted Access.  The SHMS and HMS magnets have been ramped to zero, detector HV are off (except for the SHMS scintillators), and Dave Meekins is warming up the target.

There will be no RC meeting this morning.  Meetings will resume after the shutdown.


This is a plain text email for clients that cannot display HTML.  The full logentry can be found online at https://logbooks.jlab.org/entry/3966918
-------------- next part --------------
An HTML attachment was scrubbed...
URL: <https://mailman.jlab.org/pipermail/hallc_running/attachments/20211221/e0f5e2bc/attachment.html>

More information about the Hallc_running mailing list